# Photonic entanglement enhanced multidimensional spectroscopy for probing exciton correlations: theory and applications to photosynthetic aggregates

### v1

@@ III.2 Simulations
In the following, we present the numerical simulation across three subsections. First, we deploy entangled photon pairs to explore correlations encoded in a two-exciton state by exciting them via a chosen pair of one-exciton states. Two such cases are explored, for each of which we choose a different pair of mediating one-exciton states. Second, we use a transform-limited and a frequency-chirped pulse in an attempt to recover the spectral features observed in the first case. One principal aim of this study is to recover maximal number of exciton resonances observed in the previous case by using both broadband and narrowband classical photonic sources. Third, we repeat the simulation for two different exciton nonlinearity parameters (the parameter quantifying exciton correlation within the Frenkel exciton model) and using degenerate laser pulses. This study is aimed at exploring the alteration of spectral features previously observed.
@@ III.2.1 DQC spectra with entangled photon pairs
We present the simulated two-dimensional spectra obtained using entangled photon pairs in Fig. 2. Two distinct two-exciton states, $f_{39}$ and $f_{81}$, are targeted by tuning the sum frequency $\omega_{p}$ to the respective state energies (top and bottom panels, respectively). Furthermore, the frequencies of the entangled photon pairs are chosen such that $\omega_{1}=\omega_{3}$ and $\omega_{2}=\omega_{4}$. This implies that the first pair of photons is identical to the second in the simulation. Additionally, we apply the following energy matching conditions: $\omega_{1}\approx E_{e_{05}}$ and $\omega_{2}\approx E_{e_{09}}$ for the top panel (with $\omega_{p}\approx E_{f_{39}}$), and $\omega_{1}\approx E_{e_{07}}$ and $\omega_{2}\approx E_{e_{13}}$ for the bottom panel (with $\omega_{p}\approx E_{f_{81}}$). The results presented in the left column (for both the upper and lower) serve as reference ($\tau_{p}=50.0\,\text{fs}$, $T_{\text{ent}}=10\,\text{fs}$). We intend to monitor the spectral variation with respect to the variation of temporal width of the SPDC pump ($\tau_{p}=100.0\,\text{fs}$ for the top, and $\tau_{p}=20.0\,\text{fs}$ for the bottom panel) and the entanglement time ($T_{\text{ent}}=60\,\text{fs}$), presented in the middle and right column, respectively. For each of these cases, all other parameters are kept the same as the reference.
@@
The top-left panel shows dominant resonances partitioned into five sectors along the $\tilde{\Omega}_{3}$ axis and four sectors along the $\tilde{\Omega}_{2}$ axis (where $\tilde{\Omega}=\Omega/10^{3}$ in $\text{cm}^{-1}$). These features can be identified along the $\tilde{\Omega}_{2}$ axis at $31.5$ ($\approx z_{f_{51}g}$), $30.9$ ($\approx z_{f_{49}g}$), $30.4\text{--}30.5$ ($\approx z_{f_{25}g}\text{--}z_{f_{27}g}$), and $30.2$ ($\approx z_{f_{19}g}$). Identifying the resonances along the $\tilde{\Omega}_{3}$ axis is more difficult due to interfering contributions; however, we approximately identify the peaks with the following inter-manifold resonances: $15.0$ ($\approx z_{f_{38}e_{9}}\text{--}z_{f_{38}e_{10}}$), $15.3$ ($\approx z_{f_{38}e_{8}},z_{f_{39}e_{9}}$), $15.5$ ($\approx z_{f_{38}e_{3}},z_{f_{39}e_{3}},z_{f_{40}e_{3}}$), and $15.9\text{--}16.0$ ($\approx z_{f_{41}e_{4}}$). In the top-middle panel, an increase in the temporal width produced noticeable changes by drastically reducing the spectral weights of several off-diagonal peaks. This reduction is attributed to destructive interference among the many pathways allowed by the larger bandwidth of the SPDC pump. In the top-right panel, increasing the entanglement time alters the intervals between successive exciton-photon interactions, which leads to the removal of off-diagonal spectral peaks. Consequently, only the resonance at $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(30.9,15.3)$ survives in the top-right panel, while $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(31.3,16.0)$ survives in the bottom-right panel.
@@
The bottom-left panel exhibits resonances similar to those in the top-left panel, except for the off-diagonal resonances at $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(31.5,15.5)$ and within the range $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(31.6\text{--}31.9,16.0)$ which are enhanced. An interesting case emerges when increasing the temporal width of the SPDC pump: several off-diagonal peaks survive at $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(31.5,15.3)$, $(31.5,15.5)$, $(30.7,15.5)$, and $(30.7,16.0)$. The bottom-right panel demonstrates a marked loss of resonances, with the exception of the surviving off-diagonal peak at $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(31.3,16.0)$. Overall, we note that the spectra in both panels retain their qualitative features along the $\tilde{\Omega}_{3}$ axis, even though they exhibit different prominent peaks along the $\tilde{\Omega}_{2}$ axis. In contrast to the top panel, the bottom panel exhibits reduced sensitivity to changes in the temporal width of the SPDC pump. This observation highlights that exciton correlation contributions manifest differently in these two cases.
@@ III.2.2 DQC spectra with transformed limited and chirped laser pulses
In Fig. 3, we present the two-dimensional spectra simulated using two sets of laser pulses: transform-limited and chirped laser pulses, both characterized by a Gaussian profile. The central frequencies correspond to those utilized for the entangled photon pairs, meaning that the targeted and mediating states remain consistent with those in Fig. 2. For the reference simulations, presented in the left column of both panels, we employ a transform-limited pulse with a temporal width of $\tau_{j,0}=10\,\,\text{fs}$ . We intend to explore the spectral features resulting from the variation of the temporal width ($\tau_{0}=20$ fs) and the application of an additional spectral phase (a linear chirp parameter of $750\,\text{fs}^{2}$), presented in the middle and right columns, respectively. All other parameters remain identical to those of the reference simulation.
In the left column, we observe that compared to the previous simulations in Fig. 2, the number of resolved resonances is reduced, with this reduction being more prominent along the $\Omega_{3}$ axis. Several spectral peaks exhibit significant broadening, and many carry diminished spectral weights. These features persist across all the plots. We also note that the upper-left panel exhibits a loss of a higher number of off-diagonal resonances compared to the lower one. In particular, the spectra in the upper-left, exhibit dominant resonances at $(\tilde{\Omega}_{2};\tilde{\Omega}_{3})\rightarrow(30.5;15.5),(30.7;15.5)$ which can be attributed as off-diagonal peaks. The longer temporal width of the pulses i.e., relatively narrowband case, in the upper-middle reveals that compared to the left ones, the majority of the peaks are absent at higher (at upper-middle) and lower (at lower-middle) values of $\tilde{\Omega}_{3}$. In the lower-middle, two peaks at $(\tilde{\Omega}_{2};\tilde{\Omega}_{3})\rightarrow(30.7;16.0),(30.2;16.0)$ gained spectral weights. In negatively chirped pulses, the instantaneous frequency decreases over time. In other words, the high-frequency components arrive earlier within the pulse profile. Such pulses were used to influence the discriminative step where stimulated emission to one-exciton states occurs. The inability of the parameters to explore more resonances is clearly reflected in the spectra, where we observe a few set of peaks i.e., $(\tilde{\Omega}_{2};\tilde{\Omega}_{3})\rightarrow(30.4;15.5)$ (in the upper-right) and $(\tilde{\Omega}_{2};\tilde{\Omega}_{3})\rightarrow(31.2;16.0)$ (in the lower-right) gaining significant spectral weights while others vanish. In drawing such a conclusion, we recall that the laser frequencies address the same mediating one-exciton states, making this specific combination of parameters, namely the chirp parameters and central frequencies, jointly unfavorable. It also highlights the dominant role played by two-exciton coherence generation as compared to the subsequent projection step.
@@ III.2.3 DQC spectra for a variation of the exciton nonlinearity parameters and degenerate frequencies
As described by the second term of Eq. 6, the spectral peaks and their respective weights are dictated by interfering pathways. These pathways contain two inter-manifold resonances that correspond to the energy gaps $\omega_{fe^{\prime}}$ and $\omega_{e^{\prime}g}$; the differences between these gaps determine the signal. In the Frenkel exciton description, the observed differences are largely due to the exciton nonlinearity parameters $U_{m}^{(1)}$ and $U_{mn}^{(2)}$. The exciton nonlinearities, while directly affecting the two-exciton eigenenergies, also changes the transformation matrices and, consequently, the resulting transition dipole moments. Figure 4 illustrates the influence of the variation of exciton nonlinearity parameter $U_{m}^{(1)}$ on the spectral features (left and middle columns). We also investigate the effect of degenerate frequencies of entangled photons (top-right) and classical lasers (bottom-right) to explore spectral changes caused by different $U_{m}^{(1)}$ values. Note that $U_{m}^{(1)}$ values were negative in previous sections, signifying the energetic stabilization of overtone two-exciton energies; the distribution of these values is shown in Fig. 5.
@@
The upper-left and upper-middle panels repeat the reference simulation from Fig. 2, setting the nonlinearity parameter $U_{m}^{(1)}$ to a higher (increasing them to the values depicted in Fig. 5) and lower values (lowering them to one-order of magnitude), respectively. Changing the magnitude of the $U_{m}^{(1)}$ effectively down-shifts (up-shifts) the energy gaps. It dominantly affects two-exciton eigenstates that have a higher overlap with the overtone states in the site basis. We note that in the upper-left panel, compared with the reference in Fig. 2, the peaks along $\tilde{\Omega}_{3}\approx(15.0,15.3,16.0)$ are the most affected. Specifically, their spectral weights are generally diminished, and the peaks are clustered within the range $\tilde{\Omega}_{2}\approx(30.2\text{--}31.0)$. In a similar spirit, in the lower panel, we use classical transform-limited laser with parameters identical to the reference simulation in Fig. 3. The only exception is the nonlinearity parameter, which is set to larger and smaller values in the bottom-left and bottom-middle panels, respectively. We note that in the lower-left, the peaks around $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(30.3,15.5)$ exhibit reduced spectral weight, following a trend similar to that observed for the entangled photon pairs. The previous argument remains valid: the shift in the resonances rendered the existing laser parameters unable to track the resonance frequencies. In the lower-middle, due to the reduction of exciton nonlinearity parameter by one order of magnitude, we approach a limit where two-exciton eigenenergies are approximately equal to the sum of one-exciton energies. This implies that the respective $z_{fe^{\prime}}$ resonances are of approximately equal strength to the $z_{e^{\prime}g}$. As a result, the spectra are devoid of most correlation features, except for minor residuals.
@@
The upper-right panel uses degenerate entangled photon pair frequencies to revisit the scenario presented in the top-middle panel. These frequencies are chosen such that the mediating one-exciton states become $e_{13}$. We note that two off-diagonal resonance zones, $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(31.5,15.5)$ and $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(31.5,16.0)$, become prominent, while the diagonal peaks at $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(30.7,15.5)$ are diminished. The bottom-right panel uses degenerate classical laser frequencies to revisit the scenario from the bottom-left panel. These frequencies ensure that the mediating one-exciton states remain $e_{09}$. All residual off-diagonal resonances are now significantly diminished.
@@ IV Conclusion and outlook
In this article, we demonstrated that the non-classical correlation of entangled photon pairs can be utilized to enhance the resolution of two-dimensional double quantum coherence spectra, even in the presence of dephasing. The proposed spectroscopic scheme may be realized through a combination of interferometric detection and phase-cycling techniques [48, 21, 15]. The mathematical expressions in Eq. 6 in section III.1 reveal that the nonlinear excitonic response involves three types of inter-manifold resonances: $z_{fg}$, $z_{fe}$, and $z_{eg}$. The entangled photon pairs were able to monitor and selectively amplify some of these resonances in the spectra, as shown in Fig. 2. Molecular aggregates which are composed of multiple sites, each hosting multi-level systems, the number of optical transitions are often large. When weakly coupled phonon reservoirs are present two or more dynamical resonances may involve distinct transition energies yet possess similar dephasing rates. A well characterized source of entangled photon pairs can be used to simultaneously control the photon-exciton interaction delays and combination frequencies. In turn, it may enable the discrimination between the dynamical resonances, and lead to the enhancement of the spectral resolution across either of the parametric axes, $\tilde{\Omega}_{2}$ or $\tilde{\Omega}_{3}$.
Classical transform-limited pulses, while otherwise suitable for probing ultrafast dynamics, have proved inadequate in comparable scenarios. As shown in Fig. 3 and discussed in Section III.2.2, the spectral weights associated with the two pathways could not be modulated over the full parameter regime, resulting in only a partial suppression or amplification of the allowed resonances. Appropriate spectral phase modulation is known, in certain cases, to circumvent some of the previous issues by spreading the frequency components across the pulse profile. This led us to explore the role of frequency-chirped laser pulses, as also presented in section III.2.2. We note that even though more sophisticated pulse-shaping protocols may prove useful [12, 1, 30, 35, 13, 25], the signal would still exhibit a quartic scaling with the intensity of the incoming field.
@@
The exciton response function and the resulting spectra were computed using a microscopic model of the photosynthetic complex, LHCII. The simulation accounts for both one- and two-exciton states and uses a set of realistic exciton and phonon parameters at finite temperature [45, 16]. In previous studies, ultrafast energy transfer in the one-exciton manifold was investigated spectroscopically [20, 38, 55, 2]. They often omit the role of excited-state absorption pathways, which are essential for the generation of the DQC signal.
@@
The simulation indicates that a complex interplay of exciton hopping, dephasing, and exciton nonlinearity governs the spectral weights of the signal. Within the Frenkel exciton model, the exciton nonlinearity parameter captures the effect of exciton correlations, the latter originating from Coulomb interactions. The two-exciton Hamiltonian is explicitly dependent on exciton nonlinearity parameters. The inter-manifold resonances, via the diagonalization step, carry this dependency into the expression of the signal. To assess the correlation origin of certain observed features, we varied the exciton nonlinearity parameter in Fig. 4 (and discussed in section III.2.3) and studied the resulting spectral modulation. It is demonstrated that lowering the nonlinearity parameter caused most resonances to vanish, while increasing it gave rise to a new set of resonances. We note that due to the multi-site nature of the system, a complete suppression of the signal can not be expected. The latter bolsters the capability of the DQC signal to track spectral features arising from exciton interactions and correlations across a diverse parameter regime.
@@
The Coulomb interactions between two participating exciton transitions, in general, may lead to several interesting multi-exciton phenomena such as exciton fusion, scattering, and annihilation. They await detailed spectroscopic investigation in the ultrafast regime. In the simulation, we strategically focused on two-exciton states, $f_{39}$ and $f_{81}$, lying in the lower and upper halves of the two-exciton manifold, which may be prone to exciton-exciton annihilation (EEA). Even though the spectral signatures of EEA are typically monitored via fifth-order nonlinear optical signals, lower-order optical signals, such as DQC, can already serve as an early-stage characterization tool. The DQC technique also provides a transparent protocol for the assignment of state-specific dephasing parameters involving two-exciton states. The parameter regime explored in this work provides a transparent roadmap for measurements seeking to quantify excited-state dephasing parameters using entangled photons [36]. The latter could be useful for both spectroscopy and microscopy [52, 4].
The parameters of the entangled photon pairs are determined by the properties of JSA, depicted in Fig. 6 in Appendix C, and discussed in III.1. A spectral engineering approach to the JSA would allow a consistent framework for determining the SPDC parameters [23, 51, 53]
@@
The principal outcome of this work was the ability to explore and monitor the survival of spectral signatures originating from interfering pathways. The identification of those pathways, captured in the expression containing a sum over states, and the associated photon correlation functions offers ample room for adopting open-loop control techniques. The control objectives may seek to optimize the photon correlation functions such that certain terms in the pathways are completely suppressed. It will provide an extension to the current approach, where the entangled photon parameters were microscopically informed yet intuitively estimated.
We note that applying the proposed technique to monitor the dynamics in correlated materials at the Mott limit provides an exciting opportunity [8, 17, 42, 6], which will be addressed in a subsequent communication.

### v2

@@ III.2 Simulations
In the following, we present numerical simulations across three subsections. First, we deploy entangled photon pairs to explore correlations encoded in a two-exciton state by exciting the system via a selected pair of one-exciton states. Two such cases are explored. For each of these cases we choose a different pair of mediating one-exciton states. Second, we use a transform-limited pulse and a frequency-chirped pulse to recover the spectral features observed in the first case. A principal aim of this study is to recover the maximum number of exciton resonances observed in the previous case by utilizing both broadband and narrowband classical photonic sources. Third, we repeat some of the simulations for two different exciton nonlinearity parameters and degenerate laser pulses. The exciton nonlinearity parameters quantify exciton correlation within the Frenkel exciton model. This study aims to explore the alteration of the previously observed spectral features.
@@ III.2.1 DQC spectra with entangled photon pairs
In Fig. 2, we present the simulated two-dimensional spectra obtained using entangled photon pairs. Two distinct two-exciton states, $f_{39}$ and $f_{81}$, are targeted by tuning the sum frequency $\omega_{p}$ to the respective state energies (presented in the top and bottom panels, respectively). Furthermore, the frequencies of the entangled photon pairs are selected such that $\omega_{1}=\omega_{3}$ and $\omega_{2}=\omega_{4}$, implying that the first pair of photons is identical to the second. We also apply the following energy matching conditions for selecting the mediating one-exciton states: $\omega_{1}\approx E_{e_{05}}$ and $\omega_{2}\approx E_{e_{09}}$ for the top panel (where $\omega_{p}\approx E_{f_{39}}$), and $\omega_{1}\approx E_{e_{07}}$ and $\omega_{2}\approx E_{e_{13}}$ for the bottom panel (where $\omega_{p}\approx E_{f_{81}}$). The results in the top-left and bottom-left panels utilize $\tau_{p}=50.0\,\text{fs}$, $T_{\text{ent}}=10\,\text{fs}$, which serve as the reference simulation. In the middle and right columns, we vary the temporal width of the SPDC pump, $\tau_{p}$, and the entanglement time, $T_{\text{ent}}$, respectively. For the top-middle and bottom-middle panels, we use $\tau_{p}=100.0\,\text{fs}$ and $\tau_{p}=20.0\,\text{fs}$, respectively. For both panels in the right column, we set $T_{\text{ent}}=60$ fs.
@@
The top-left panel shows a series of peaks which can be partitioned into five sectors along the $\tilde{\Omega}_{3}$ axis and four sectors along the $\tilde{\Omega}_{2}$ axis (where $\tilde{\Omega}=\Omega/10^{3}$ in $\text{cm}^{-1}$). Along the $\tilde{\Omega}_{2}$ axis, these spectral features can be identified can be identified as $31.5$ ($\approx z_{f_{51}g}$), $30.9$ ($\approx z_{f_{49}g}$), $30.4\text{--}30.5$ ($\approx z_{f_{25}g}\text{--}z_{f_{27}g}$), and $30.2$ ($\approx z_{f_{19}g}$). Identifying the peaks along the $\tilde{\Omega}_{3}$ axis is more cumbersome due to interfering contributions. However, we approximately identify the peaks with the following inter-manifold resonances: $15.0$ ($\approx z_{f_{38}e_{9}}\text{--}z_{f_{38}e_{10}}$), $15.3$ ($\approx z_{f_{38}e_{8}},z_{f_{39}e_{9}}$), $15.5$ ($\approx z_{f_{38}e_{3}},z_{f_{39}e_{3}},z_{f_{40}e_{3}}$), and $15.9\text{--}16.0$ ($\approx z_{f_{41}e_{4}}$). In the top-middle panel, changing the temporal width produced several noticeable changes; the spectral weights of several off-diagonal peaks are drastically reduced. The reduction is attributed to lowering of the spectral bandwidth which allows fewer pathways. Increasing the entanglement time alters the intervals between successive exciton-photon interactions. In the top-right panel, it leads to the vanishing of off-diagonal peaks. Only the peaks at $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(30.9,15.3)$ survives in the top-right panel, while $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(31.3,16.0)$ survives in the bottom-right panel.
@@
The bottom-left panel exhibits peaks similar to those in the top-left panel, except for the off-diagonal peaks at $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(31.5,15.5)$ and within the range $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(31.6\text{--}31.9,16.0)$ which are now pronounced. An interesting case emerges when increasing the temporal width of the SPDC pump: several off-diagonal peaks survive at $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(31.5,15.3)$, $(31.5,15.5)$, $(30.7,15.5)$, and $(30.7,16.0)$. The larger spectral bandwidth allows some new peaks while retaining most of the earlier ones. This can be explained by destructive interference among the pathways allowed by the larger bandwidth. The bottom-right panel demonstrates a marked loss in the number of spectral features, with the exception of a surviving off-diagonal peak at $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(31.3,16.0)$. Overall, we note that the spectra in both panels retain their qualitative features along the $\tilde{\Omega}_{3}$ axis, even though they exhibit differences along the $\tilde{\Omega}_{2}$ axis. In contrast to the top panel, the bottom panel exhibits reduced sensitivity to changes in the temporal width of the SPDC pump. This observation highlights that exciton correlation contributions manifest differently for the targeted $|f\rangle$ states.
@@ III.2.2 DQC spectra with transformed limited and chirped laser pulses
In Fig. 3, we present the two-dimensional spectra simulated using two sets of classical laser pulses: transform-limited and chirped laser, both characterized by a Gaussian profile. The central frequencies correspond to those utilized for the entangled photon pairs, meaning that the targeted and mediating states remain consistent with those in Fig. 2. For the reference simulations, presented in the left column of both panels, we employ a transform-limited pulse with a temporal width of $\tau_{j,0}=10\,\,\text{fs}$ . We explore the spectral features resulting from the variation of the temporal width, by setting $\tau_{j,0}=20$ fs, and the application of an additional spectral phase, by using a linear chirp parameter of 750-750 fs2, in the middle and right columns, respectively.
In the left column, we observe that compared to the previous simulations in Fig. 2, the number of spectrally resolved resonances is reduced, particularly along the $\Omega_{3}$ axis. Several spectral peaks exhibit marked broadening, and many carry diminished spectral weights. These features persist across all the plots. We also note that the top-left panel exhibits a loss of more number of off-diagonal peaks compared with the bottom-left panel. In particular, the spectra in the top-left exhibit dominant off-diagonal peaks at $(\tilde{\Omega}_{2};\tilde{\Omega}_{3})\rightarrow(30.5;15.5),(30.7;15.5)$. The longer temporal width of the pulse indicate relatively narrow spectral width. The middle panel reveals that compared to the left panels, the majority of the peaks in the top- and bottom-middle panels are absent at higher and lower values of $\tilde{\Omega}_{3}$, respectively. In the bottom-middle panel, two peaks at $(\tilde{\Omega}_{2};\tilde{\Omega}_{3})\rightarrow(30.7;16.0),(30.2;16.0)$ gained spectral weights. In negatively chirped pulses, the instantaneous frequency decreases over time. In other words, the high-frequency components arrive earlier within the pulse profile. Such pulses can be used to influence the discriminative stage of the DQC signal, where stimulated emission involving one-exciton resonances occurs. The limited capacity of the chirp parameters to resolve multiple resonances is reflected in the spectra. We observe a few peaks at $(\tilde{\Omega}_{2};\tilde{\Omega}_{3})\rightarrow(30.4;15.5)$ in the top-right panel and $(\tilde{\Omega}_{2};\tilde{\Omega}_{3})\rightarrow(31.2;16.0)$ in the bottom-right panel gaining significant spectral weights while others vanish. In drawing this conclusion, we note that since the classical laser frequencies address the same mediating one-exciton states, the specific combination of the chirp parameters and central frequencies is unfavorable. This result also highlights the dominant role played by two-exciton coherence generation stage.
@@ III.2.3 DQC spectra for a variation of the exciton nonlinearity parameters and degenerate frequencies
As described by the second term of Eq. 6, the spectral peaks and their respective weights are dictated by interference between two pathways. These pathways contain inter-manifold resonances that correspond to the energy gaps $\omega_{fe^{\prime}}$ and $\omega_{e^{\prime}g}$; the differences between these resonances determine the overall magnitude of the signal. In the Frenkel exciton description, these differences are largely due to the exciton nonlinearity parameters $U_{m}^{(1)}$ and $U_{mn}^{(2)}$. The exciton nonlinearities, while directly affecting the two-exciton eigenenergies, also changes the transformation matrices and, consequently, the transition dipole moments. Figure 4 illustrates the influence of the variation of exciton nonlinearity parameter $U_{m}^{(1)}$ on the spectra, presented in the left and middle columns. Note that $U_{m}^{(1)}$ values were negative in previous sections, signifying the energetic stabilization of overtone two-exciton energies; the distribution of these values is shown in Fig. 5. We also investigate the effect of degenerate entangled photon pairs and degenerate classical lasers on the spectra, presented in the top- and bottom-right panels, respectively.
@@
The top-left and top-middle panels repeat the reference simulation from Fig. 2, but the nonlinearity parameters $U_{m}^{(1)}$ are set to higher values, respectively. Specifically, these parameters are increased to the values depicted in Fig. 5. Setting the magnitude of the $U_{m}^{(1)}$ higher effectively shifts the energy gaps. This change dominantly affects two-exciton eigenstates that have a high overlap with the overtone states in the site basis. We note that in the top-left panel, compared with the reference in Fig. 2, the peaks along $\tilde{\Omega}_{3}\approx(15.0,15.3,16.0)$ are the most affected. In general, their spectral weights are diminished, and the peaks are clustered within the range $\tilde{\Omega}_{2}\approx(30.2\text{--}31.0)$.
In a similar spirit, in the lower panels, we use classical transform-limited laser pulses with parameters identical to those of the reference simulation in Fig. 3. The nonlinearity parameters are set to higher and lower values in the bottom-left and bottom-middle panels, respectively. The latter makes these values in the bottom-middle panel negligibly small. We note that in the bottom-left panel, the peaks around $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(30.3,15.5)$ exhibit reduced spectral weight, following a trend similar to that observed for the entangled photon pairs. The previous argument remains valid: the shift in the resonances rendered the existing laser parameters unable to track the resonance frequencies. In the bottom-middle panel, due to the reduction of the exciton nonlinearity parameter by one order of magnitude, we approach a limit where two-exciton eigenenergies are approximately equal to the sum of one-exciton energies. This implies that the respective $z_{fe^{\prime}}$ resonances are of approximately equal strength to the $z_{e^{\prime}g}$. Consequently, the spectra are devoid of most correlation features, except for minor residuals.
@@
The top-right panel utilizes degenerate entangled photon pairs to revisit the scenario presented in the top-middle panel. These frequencies are selected so that the mediating one-exciton states become $e_{13}$. We note that two off-diagonal zones around $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(31.5,15.5)$ and $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(31.5,16.0)$, become prominent, while the diagonal peaks at $(\tilde{\Omega}_{2},\tilde{\Omega}_{3})\approx(30.7,15.5)$ are diminished. The bottom-right panel utilizes degenerate classical laser pulses to revisit the scenario from the bottom-left panel. Here, we choose the central frequencies to ensure that the mediating one-exciton states remain $e_{09}$. All residual off-diagonal resonances are now markedly diminished.
@@ IV Conclusion and outlook
In this article, we demonstrated that the non-classical correlation of entangled photon pairs can be utilized to enhance the resolution of two-dimensional double quantum coherence spectra, even in the presence of dephasing. The proposed scheme can be realized by combining phase-cycling techniques with interferometric detection [48, 21, 15]. The mathematical expressions in Eq. 6 in section III.1 reveal that the nonlinear excitonic response involves three types of inter-manifold resonances: $z_{fg}$, $z_{fe}$, and $z_{eg}$. The entangled photon pairs were able to monitor and selectively amplify some of these resonances in the spectra, as shown in Fig. 2. Molecular aggregates which are composed of multiple sites, each hosting multiple energy levels, the number of plausible optical transitions are often large. It leads to a large number of dynamical resonances; they may appear congested in the spectra. Furthermore, when weakly coupled phonon modes act as reservoirs, two or more dynamical resonances may involve distinct transition energies yet possess similar dephasing rates; they appear coalesced in the spectra. A well characterized source of entangled photon pairs can simultaneously control the delays between successive exciton-photon interactions and the combination frequencies of the photons. This, in turn, may enable the discrimination between the dynamical resonances, leading to the enhancement of the spectral resolution across either of the parametric axes, $\tilde{\Omega}_{2}$ or $\tilde{\Omega}_{3}$. Classical transform-limited pulses, while otherwise suitable for probing ultrafast dynamics, have proved to be inadequate in the investigated scenarios. As shown in Fig. 3 and discussed in Section III.2.2, the spectral weights associated with the two pathways could not be modulated over the full parameter regime, resulting in only a partial suppression or amplification of the resonances. In certain cases, appropriate spectral phase modulation is known to circumvent some of the previous issues by spreading the frequency components across the pulse profile. This led us to explore the role of frequency-chirped laser pulses, discussed in the Section III.2.2. We note that even though a more sophisticated pulse-shaping protocols can be used [12, 1, 30, 35, 13, 25], the signal would still exhibit a quartic scaling with the intensity of the incoming field.
@@
The exciton nonlinear response and the resulting two-dimensional spectra were computed using a microscopic model of the photosynthetic complex, LHCII. The simulation indicates that a complex interplay of exciton hopping, dephasing, and exciton nonlinearity governs the spectral weights of the signal. In previous studies, ultrafast energy transfer in the one-exciton manifold was spectroscopically investigated [20, 38, 55, 2]. These studies often omit the role of excited-state absorption pathways, which are essential for the generation of the DQC signal. The simulation uses a set of realistic exciton and phonon parameters at finite temperature, parameters of which were taken from [45, 16, 14].
@@
Within the Frenkel exciton model, the exciton nonlinearity parameter captures the effect of exciton correlations, the latter originating from Coulomb interactions. The simulation accounts for both one- and two-exciton states; the two-exciton Hamiltonian is explicitly dependent on exciton nonlinearity parameters. The inter-manifold resonances, via the diagonalization step, carry this dependency into the expression of the signal. To examine the correlation origin of some of the observed features, we varied the exciton nonlinearity parameter and studied the resulting spectral modulation. It is discussed in Section III.2.3. It is demonstrated that lowering the nonlinearity parameter caused most resonances to vanish, while increasing it gave rise to a new set of resonances. The strategy bolsters the capability of the DQC signal to track spectral features arising from exciton interactions and correlations across a diverse parameter regime. We note that due to the presence of multiple levels in two different manifolds, a complete suppression of the signal can not be expected.
@@
The Coulomb interactions involving two-exciton transitions also lead to several interesting phenomena, such as exciton fusion, scattering, and annihilation. These processes await a detailed spectroscopic investigation in the ultrafast regime. In our simulations, we strategically focus on the two-exciton states $f_{39}$ and $f_{81}$, which lie in the lower and upper halves of the two-exciton manifold and may be prone to exciton-exciton annihilation (EEA). Even though the spectral signatures of EEA are typically monitored via fifth-order nonlinear optical signals, lower-order signals, such as DQC, can serve as an early-stage characterization tool. The DQC technique also provides a protocol for assigning state-specific dephasing parameters for the involved two-exciton states. The parameter regime explored in this work provides a transparent guide for measurements seeking to quantify excited-state dephasing parameters using entangled photons [36]. These measurements would be useful for both spectroscopy and microscopy [52, 4].
@@
The parameters of the entangled photon pairs are determined by the properties of the Joint Spectral Amplitude (JSA), depicted in Fig. 6 in Appendix C and discussed in Section III.1. A spectral engineering approach to the JSA would offer a consistent framework for determining the SPDC parameters [23, 51, 53]. The principal outcome of this work is the ability to explore and monitor the survival of spectral signatures originating from interfering pathways. The pathways in the signal expression appear in a convolutional form containing exciton and photon correlation functions; it offers ample room for adopting open-loop control techniques. Control objectives may seek to optimize the photon correlation functions such that specific terms in the sum-over-states expression are completely suppressed. It will provide an extension to the approach adopted in this work, where the entangled photon parameters were microscopically informed yet intuitively estimated. We note that applying the proposed technique to monitor the dynamics in correlated materials at the Mott limit provides an exciting opportunity [8, 17, 42, 6], which will be addressed in a subsequent communication.